OpenStudy (anonymous):

From 1960 to 1970, the consumer price index (CPI) increased from 29.6 to 48.2. If a dozen donuts cost $0.89 in 1960 and the price of donuts increased at the same rate as the CPI from 1960 to 1970, approximately how much did a dozen donuts cost in 1970?

OpenStudy (anonymous):

A. $1.63 B. $1.49 C. $1.45 D. $1.79

OpenStudy (ranga):

x / 0.89 = 48.2 / 29.6 Solve for x.

OpenStudy (anonymous):

still dnt get it

OpenStudy (ranga):

x = (0.89)(48.2)/29.6 = $1.45

OpenStudy (anonymous):

ooo i see thankss

OpenStudy (ranga):

yw